# Quantum Algorithms for the Pathwise Lasso

### v1

@@ 6      $t\leftarrow t+1$
$\geq\Big{(}1-\frac{\epsilon}{2}\Big{)}\widetilde{\lambda}_{t+1}^{\rm join}-\frac{\epsilon\widetilde{\lambda}^{\rm join}_{t+1}}{2}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}\|_{1}$
$\geq(1-\epsilon)\widetilde{\lambda}_{t+1}^{\rm join},$
@@
where we used that $\big{|}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda^{\rm join}_{t+1}))|=\lambda_{t+1}^{\rm join}$ by the definition of the joining time $\lambda_{t+1}^{\rm join}$, that $\widetilde{\bm{\beta}}(\lambda_{t})$ satisfies the $\mathsf{KKT}_{\lambda_{t}}(\epsilon)$ condition at $\lambda_{t}$, i.e., $\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}\leq\lambda_{t}$, and finally that $\max_{j\in\mathcal{A}^{c}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{j}\|_{1}\leq 1$. ∎
@@
After asserting the correctness of Algorithm 3, we now analyse its complexity. Specifically, we show how to sample the joining variable $\widetilde{i}_{t+1}^{\rm join}$ from the set in Eq. (13) by combining the approximate quantum minimum-finding subroutine from Chen and de Wolf [CdW23] (Fact 11) and the unitary map that approximates (parts of) the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ (Lemma 13). Our final complexity depends on the mutual incoherence between different column subspaces and the overlap between $\mathbf{y}$ and different column subspaces.
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{y}\in\mathbb{R}^{n}$. Assume $\mathbf{X}$ is stored in a QROM and we have access to QRAMs of size $O(n)$. Let $\delta\in(0,1)$, $\epsilon\in(0,2)$, and $T\in\mathbb{N}$. Let $\alpha,\gamma\in(0,1]$ be such that
[…]
@@
$O\left(\frac{\gamma^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}}{\alpha\epsilon}\sqrt{|\mathcal{I}|}\log^{2}(T/\delta)\operatorname{poly}\log(nd)+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq(1-\epsilon/2)\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\}$ (14)
@@
$\widetilde{\lambda}_{t+1}^{\rm join}=(1-\epsilon/2)^{-1}\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A}).$
Regarding the time complexity of the algorithm, the most expensive steps are computing $\mathbf{X}_{\mathcal{A}}^{+}$ in time $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$, inputting $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into KP-trees in time $O(n\log{n})$, and sampling $\widetilde{i}_{t+1}^{\rm join}$. We now show how to obtain the joining variable $\widetilde{i}_{t+1}^{\rm join}$, i.e., how it can be sampled from the set in Eq. (14).
@@
in time $O(\epsilon_{1}^{-1}\log(1/\delta_{0})\operatorname{poly}\log(nd))$ and $O(\epsilon_{2}^{-1}\log(1/\delta_{0})\operatorname{poly}\log(nd))$, respectively. Then, by using Lemma 8, we can construct the desired map $|i\rangle|\bar{0}\rangle\mapsto|i\rangle|R_{i}\rangle$ as in Eq. (15) with
@@
$\leq\frac{1}{\lambda_{t}}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{\infty}\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}$
$\leq\sqrt{n}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}$
$\leq\sqrt{n}\|\mathbf{X}^{+}\|_{2},$
@@
$\geq 1-\frac{1}{\lambda_{t}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{i}\|_{1}\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}$
$\geq\alpha,$
where we used Hölder’s inequality and again that $\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}\leq\lambda_{t}$. The above inequalities are sufficient to bound the second term in Eq. (16). Regarding the first term,
@@
$\leq\epsilon_{1}\alpha^{-1}\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}$
$=\epsilon_{1}\alpha^{-1}\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}\frac{\lambda_{t+1}^{\rm join}}{\max_{i\in\mathcal{I}}\left|\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}\right|}$
$\leq 2\epsilon_{1}\alpha^{-1}\lambda_{t+1}^{\rm join}\frac{\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}}{\max_{i\in\mathcal{I}}|\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})|}$
$\leq 2\epsilon_{1}\alpha^{-1}\gamma^{-1}\lambda_{t+1}^{\rm join},$
using that $\max_{i\in\mathcal{I}}|\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})|=\|\mathbf{X}_{\mathcal{I}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})\|_{\infty}=\|\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})\|_{\infty}\geq\gamma\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}$ and that $\alpha\leq|{\pm}1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}|\leq 2$. All the above leads to
$\epsilon_{0}\leq\lambda^{\rm join}_{t+1}\big{(}2\epsilon_{1}\alpha^{-1}\gamma^{-1}+\epsilon_{2}\alpha^{-1}\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}\big{)}.$
By taking (notice that $\epsilon\in(0,2)$ guarantees that $\epsilon_{2}$ is within the required range from Lemma 8)
$\epsilon_{1}=\frac{\alpha\gamma\epsilon}{8}\quad\text{and}\quad\epsilon_{2}=\frac{\alpha\epsilon}{4\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}},$
then $\epsilon_{0}\leq\epsilon\lambda^{\rm join}_{t+1}/4$ in time $O\Big{(}\frac{\gamma^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}}{\alpha\epsilon}\log(1/\delta_{0})\operatorname{poly}\log(nd)\Big{)}$. Finally, we apply the approximate quantum minimum-finding algorithm (Fact 11) to obtain an index $\widetilde{i}_{t+1}^{\rm join}\in\mathcal{I}$ such that
$\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})\geq\max_{i\in\mathcal{I}}\{\Lambda^{\rm join}_{i}(\mathcal{A})\}-2\epsilon_{0}\geq(1-\epsilon/2)\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\},$
i.e., $\widetilde{i}_{t+1}^{\rm join}$ belongs to the set from Eq. (14) with probability at least $1-\delta/T$ as promised. The total complexity is $\widetilde{O}(\sqrt{|\mathcal{I}|}\log(T/\delta))$ times the complexity of constructing the mapping $|i\rangle|\bar{0}\rangle\mapsto|i\rangle|R_{i}\rangle$. The final success probability follows from a union bound. ∎
If $\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}\|_{2}$ is constant and the parameters $\alpha,\gamma\in(0,1]$ are bounded away from 00, then the complexity of Algorithm 3 is $\widetilde{O}(\sqrt{n|\mathcal{I}|}/\epsilon+n|\mathcal{A}|+|\mathcal{A}|^{2})$ per iteration. If $|\mathcal{A}|=O(n)$, then we obtain a fully quadratic advantage $\widetilde{O}(\sqrt{nd})$ over the classical LARS algorithm.
@@ 4.3 Approximate classical LARS algorithm
By using virtually the same techniques and results behind the approximate quantum LARS algorithm, it is possible to devise an analogous approximate classical LARS algorithm. The idea is again to approximately compute the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$, but by using the classical sampling procedure from Lemma 14 instead. For well-behaved design matrices, the complexity per iteration is $\widetilde{O}(n|\mathcal{I}|/\epsilon^{2}+n|\mathcal{A}|+|\mathcal{A}|^{2})$. The approximate classical LARS algorithm for the pathwise Lasso is shown in Algorithm 4.
@@
1 Input: Vector $\mathbf{y}\in\mathbb{R}^{n}$, matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$, $\delta\in(0,1)$, $\epsilon\in(0,2)$, and $T\in\mathbb{N}$ Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\widetilde{\bm{\beta}}(\lambda_{0})=0$, $t=0$ while $\mathcal{I}\neq\varnothing$, $t\leq T$ do
       $\bm{\eta}_{\mathcal{A}}\leftarrow\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))$ $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$   // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
       $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$   // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
2       Input $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into classical-samplable memories Define $\Lambda_{i}^{\rm join}\triangleq\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and $\Lambda_{i}^{\rm cross}\triangleq\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]$ $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ Compute $\{r_{i}\}_{i\in\mathcal{I}}$ such that $|r_{i}-\Lambda_{i}^{\rm join}|\leq\frac{\epsilon\lambda_{t+1}^{\rm join}}{2}$ with failure probability $\frac{\delta}{T|\mathcal{I}|}$ (Lemma 14) $\widetilde{i}_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{r_{i}\}$ and $\widetilde{\lambda}_{t+1}^{\rm join}\leftarrow(1-\epsilon/2)^{-1}\max_{i\in\mathcal{I}}\{r_{i}\}$ $\lambda_{t+1}\leftarrow\min\{\lambda_{t},\max\{\lambda_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}\}$ $\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$ if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then

### v2

@@ 24      $t\leftarrow t+1$
$\geq\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}\widetilde{\lambda}_{t+1}^{\rm join}-\frac{\epsilon\widetilde{\lambda}^{\rm join}_{t+1}}{1+\alpha_{\mathcal{A}}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\widetilde{i}_{t+1}^{\rm join}}\|_{1}$ ($\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}\geq\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}\widetilde{\lambda}_{t+1}^{\rm join}$)
$\geq(1-\epsilon)\widetilde{\lambda}_{t+1}^{\rm join},$ ($\max_{j\in\mathcal{A}^{c}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{j}\|_{1}\leq\alpha_{\mathcal{A}}$)
@@
where we used that $\widetilde{\bm{\beta}}(\lambda_{t})$ satisfies the $\mathsf{KKT}_{\lambda_{t}}(\epsilon)$ condition at $\lambda_{t}$, $\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}\leq\lambda_{t}$. ∎
@@
After asserting the correctness of Algorithm 3, we now analyse its complexity. Specifically, we show how to sample the joining variable $\widetilde{i}_{t+1}^{\rm join}$ from the set in Equation 11 by combining the approximate quantum minimum-finding subroutine from Chen and de Wolf [23] (Fact 14) and the unitary map that approximates (parts of) the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ (Lemma 16). Our final complexity depends on the mutual incoherence between different column subspaces and the overlap between $\mathbf{y}$ and different column subspaces.
@@
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ and $\mathbf{y}\in\mathbb{R}^{n}$. Assume $\mathbf{X}$ is stored in a QROM and we have access to QRAMs of size $O(n)$. Let $\delta\in(0,1)$, $\epsilon>0$, and $T\in\mathbb{N}$. For $\mathcal{A}\subseteq[d]$ of size $|\mathcal{A}|\leq T$, assume there are $\alpha_{\mathcal{A}},\gamma_{\mathcal{A}}\in(0,1)$ such that
$\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\leq\alpha_{\mathcal{A}}\quad\text{and}\quad\frac{\|\mathbf{X}^{\top}(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{\infty}}{\|\mathbf{X}\|_{\max}\|(\mathbf{I}-\mathbf{X}_{\mathcal{A}}\mathbf{X}_{\mathcal{A}}^{+})\mathbf{y}\|_{1}}\geq\gamma_{\mathcal{A}}.$
@@
$O\left(\frac{\gamma^{-1}_{\mathcal{A}}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\sqrt{|\mathcal{I}|}\log^{2}(T/\delta)\operatorname{poly}\log(nd)+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
@@
$\left\{j\in\mathcal{I}:\Lambda_{j}^{\rm join}(\mathcal{A})\geq\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\}\right\}$ (12)
@@
$\widetilde{\lambda}_{t+1}^{\rm join}=\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}^{-1}\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A}).$
Regarding the time complexity of the algorithm, the most expensive steps are computing $\mathbf{X}_{\mathcal{A}}^{+}$ in time $O(n|\mathcal{A}|+|\mathcal{A}|^{2})$, inputting $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into KP-trees in time $O(n\log{n})$, and sampling $\widetilde{i}_{t+1}^{\rm join}$. We now show how to obtain the joining variable $\widetilde{i}_{t+1}^{\rm join}$, i.e., how to sample it from the set in Equation 12. In the following, fix $\mathcal{A}$ and $\mathcal{I}$.
@@
in time $O(\epsilon_{1}^{-1}\log(1/\delta_{0})\operatorname{poly}\log(nd))$ and $O(\epsilon_{2}^{-1}\log(1/\delta_{0})\operatorname{poly}\log(nd))$, respectively. Then, by using Lemma 10, we can construct the desired map $|i\rangle|\bar{0}\rangle\mapsto|i\rangle|R_{i}\rangle$ as in Equation 13 with
@@
$\leq\frac{1}{\lambda_{t}}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{\infty}\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}$ (Hölder’s inequality)
$\leq\|\mathbf{X}_{\mathcal{A}}^{+}\|_{\infty}$ ($\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}\leq\lambda_{t}$)
$\leq\sqrt{n}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2},$ ($\|\mathbf{X}_{\mathcal{A}}^{+}\|_{\infty}\leq\sqrt{n}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}$)
@@
$\geq 1-\frac{1}{\lambda_{t}}\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{i}\|_{1}\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}$ (Hölder’s inequality)
$\geq 1-\alpha_{\mathcal{A}},$ ($\|\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))\|_{\infty}\leq\lambda_{t}$ and $\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{i}\|_{1}\leq\alpha_{\mathcal{A}}$)
The above inequalities are sufficient to bound the second term in Equation 14. Regarding the first term,
@@
$\leq\frac{\epsilon_{1}}{1-\alpha_{\mathcal{A}}}\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}$
$=\frac{\epsilon_{1}}{1-\alpha_{\mathcal{A}}}\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}\frac{\lambda_{t+1}^{\rm join}}{\max_{i\in\mathcal{I}}\left|\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}\right|}$
$\leq\frac{(1+\alpha_{\mathcal{A}})\epsilon_{1}\lambda_{t+1}^{\rm join}}{1-\alpha_{\mathcal{A}}}\frac{\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}}{\max_{i\in\mathcal{I}}|\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})|}$
$\leq\frac{(1+\alpha_{\mathcal{A}})\epsilon_{1}\gamma_{\mathcal{A}}^{-1}\lambda_{t+1}^{\rm join}}{1-\alpha_{\mathcal{A}}},$
using that $\max_{i\in\mathcal{I}}|\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})|=\|\mathbf{X}_{\mathcal{I}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})\|_{\infty}=\|\mathbf{X}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})\|_{\infty}\geq\gamma_{\mathcal{A}}\|\mathbf{X}\|_{\max}\|\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}\|_{1}$ and that $1-\alpha_{\mathcal{A}}\leq|{\pm}1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}|\leq 1+\alpha_{\mathcal{A}}$. All the above leads to
$\epsilon_{0}\leq\frac{\epsilon_{1}(1+\alpha_{\mathcal{A}})\gamma_{\mathcal{A}}^{-1}+\epsilon_{2}\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}}{1-\alpha_{\mathcal{A}}}\lambda^{\rm join}_{t+1}.$
By taking
$\epsilon_{1}=\frac{\epsilon(1-\alpha_{\mathcal{A}})\gamma_{\mathcal{A}}}{2(1+\alpha_{\mathcal{A}})^{2}}\quad\text{and}\quad\epsilon_{2}=\frac{\epsilon(1-\alpha_{\mathcal{A}})}{2(1+\alpha_{\mathcal{A}})\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}_{\mathcal{A}}^{+}\|_{2}},$
then $\epsilon_{0}\leq\frac{\epsilon}{2(1+\alpha_{\mathcal{A}})}\lambda^{\rm join}_{t+1}$ in time $O\Big{(}\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\log(1/\delta_{0})\operatorname{poly}\log(nd)\Big{)}$. Finally, we apply the approximate quantum minimum-finding algorithm (Fact 14) to obtain an index $\widetilde{i}_{t+1}^{\rm join}\in\mathcal{I}$ such that
$\Lambda^{\rm join}_{\widetilde{i}_{t+1}^{\rm join}}(\mathcal{A})\geq\max_{i\in\mathcal{I}}\{\Lambda^{\rm join}_{i}(\mathcal{A})\}-2\epsilon_{0}\geq\Big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\Big{)}\max_{i\in\mathcal{I}}\{\Lambda_{i}^{\rm join}(\mathcal{A})\},$
i.e., $\widetilde{i}_{t+1}^{\rm join}$ belongs to the set from Equation 12 with probability at least $1-\delta/T$ as promised. The total complexity is $\widetilde{O}(\sqrt{|\mathcal{I}|}\log(T/\delta))$ times the complexity of constructing the mapping $|i\rangle|\bar{0}\rangle\mapsto|i\rangle|R_{i}\rangle$. The final success probability follows from a union bound. ∎
[…]
@@ 4.3 Approximate classical LARS algorithm
By using similar techniques and results behind the approximate quantum LARS algorithm, it is possible to devise an analogous approximate classical LARS algorithm, which can be seen as a dequantised counterpart to our quantum algorithms. The idea is again to approximately compute the joining times $\Lambda_{i}^{\rm join}(\mathcal{A})=\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$, but by using the classical sampling procedure from Lemma 17 instead. For well-behaved design matrices, the complexity per iteration is $\widetilde{O}(n|\mathcal{I}|/\epsilon^{2}+n|\mathcal{A}|+|\mathcal{A}|^{2})$. The approximate classical LARS algorithm for the pathwise Lasso is shown in Algorithm 4.
1
Input: $T\in\mathbb{N}$, $\delta\in(0,1)$, $\epsilon>0$, $\mathbf{y}\in\mathbb{R}^{n}$, $\mathbf{X}\in\mathbb{R}^{n\times d}$, and $\{\alpha_{\mathcal{A}}\geq\|\mathbf{X}_{\mathcal{A}}^{+}\mathbf{X}_{\mathcal{A}^{c}}\|_{1}\}_{\mathcal{A}\subseteq[d]:|\mathcal{A}|\leq T}$
2
Output: Regularisation path $\widetilde{\mathcal{P}}$ with $T$ kinks and error $\lambda\epsilon\|\widetilde{\bm{\beta}}(\lambda)\|_{1}$ with probability $\geq 1-\delta$
3
@@
4Initialise $\mathcal{A}=\operatorname{argmax}_{j\in[d]}|\mathbf{X}_{j}^{\top}\mathbf{y}|$, $\mathcal{I}=[d]\setminus\mathcal{A}$, $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$, $\widetilde{\bm{\beta}}(\lambda_{0})=\mathbf{0}$, $t=0$
5while $\mathcal{I}\neq\varnothing$, $t\leq T$ do
6       $\bm{\eta}_{\mathcal{A}}\leftarrow\frac{1}{\lambda_{t}}\mathbf{X}_{\mathcal{A}}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t}))$
      $\bm{\mu}\leftarrow\mathbf{X}_{\mathcal{A}}^{+}\mathbf{y}$
        // $\bm{\mu}\in\mathbb{R}^{|\mathcal{A}|}$
7
      $\bm{\theta}\leftarrow(\mathbf{X}_{\mathcal{A}}^{\top}\mathbf{X}_{\mathcal{A}})^{+}\bm{\eta}_{\mathcal{A}}$
        // $\bm{\theta}\in\mathbb{R}^{|\mathcal{A}|}$
8
9      Input $\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu}$ and $\mathbf{X}_{\mathcal{A}}\bm{\theta}$ into classical-samplable memories
10      Define $\Lambda_{i}^{\rm join}\triangleq\frac{\mathbf{X}_{i}^{\top}(\mathbf{y}-\mathbf{X}_{\mathcal{A}}\bm{\mu})}{\pm 1-\mathbf{X}_{i}^{\top}\mathbf{X}_{\mathcal{A}}\bm{\theta}}$ and $\Lambda_{i}^{\rm cross}\triangleq\frac{\mu_{i}}{\theta_{i}}\cdot\mathbf{1}\left[\frac{\mu_{i}}{\theta_{i}}\leq\lambda_{t}\right]$
11      $i_{t+1}^{\rm cross}\leftarrow\operatorname{argmax}_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$ and $\lambda_{t+1}^{\rm cross}\leftarrow\max_{i\in\mathcal{A}}\{\Lambda_{i}^{\rm cross}\}$
12      Compute $\{r_{i}\}_{i\in\mathcal{I}}$ such that $|r_{i}-\Lambda_{i}^{\rm join}|\leq\frac{\epsilon\lambda_{t+1}^{\rm join}}{1+\alpha_{\mathcal{A}}}$ with failure probability $\frac{\delta}{T|\mathcal{I}|}$ (Lemma 17)
13      $\widetilde{i}_{t+1}^{\rm join}\leftarrow\operatorname{argmax}_{i\in\mathcal{I}}\{r_{i}\}$ and $\widetilde{\lambda}_{t+1}^{\rm join}\leftarrow\big{(}1-\frac{\epsilon}{1+\alpha_{\mathcal{A}}}\big{)}^{-1}\max_{i\in\mathcal{I}}\{r_{i}\}$
14      $\lambda_{t+1}\leftarrow\min\{\lambda_{t},\max\{\lambda_{t+1}^{\rm cross},\widetilde{\lambda}_{t+1}^{\rm join}\}\}$
15
16      $\widetilde{\bm{\beta}}_{\mathcal{A}}(\lambda_{t+1})\leftarrow\bm{\mu}-\lambda_{t+1}\bm{\theta}$
17
18      if $\lambda_{t+1}=\lambda_{t+1}^{\rm cross}$ then